The golf ball manufacturer has developed a profit model that depends on the number x of golf balls 

sold per month(measured in thousands),and the number of hours per month of advertising y, according 

to the function

𝑧 = 𝑓(𝑥, 𝑦) = 48𝑥 + 96𝑦 − 𝑥

2 − 2𝑥𝑦 − 9𝑦

2

,

Where z is measured in thousands of dollars .The budgetary constraint function relating the cost of the 

production of thousands golf balls and advertising units is given by 

20𝑥 + 4𝑦 = 216,

Find the values of x and y that maximize profit, and find the maximum profit


1
Expert's answer
2021-05-19T16:49:38-0400

Given,

"f(x,y)=48x+96y-x^2-2xy-9y^2"


Constraints : "20x+4y=216=\phi(x,y)~~~~~~~~~-(1)"


By suing Lagrange's method-


"F(x,)=f(x,y)+\lambda \phi(x,y)"


    "=48x+96y-x^2-2xy-9y^2+\lambda(20x+4y-216)"


  

Now, "F'_x(x,y)=48-2x-2y+20\lambda=0~~~~~~~-(2)"


  "F'_y(X,y)=-96-2x-18y+4\lambda=0~~~~~~~" -(3)


From equation 2 we have ,"\lambda=\dfrac{2x+2y-48}{20}"


    From eqn.(3) "\lambda=\dfrac{2x+18y-96}{4}"


Equating the above 2 values of "\lambda-"


"\dfrac{2x+2y-48}{205}=\dfrac{2x+18y-96}{4}

\\[9pt]

2(x+y-24)=10(x+9y-48)
\\[9pt]


x+11y-54=0~~~~~-(4)"


Solving equation (1) and (4) and we get-


x=10,y=4


Hence, Maximum Profit


"Z=f(10,4)=48(10)+96(4)-10^2-2(10)(4)-9(4)^2
\\
 =480+384-100-80-144\\

 =540"